Calculate the molar mass of the following substances.
       (a) Ethyne, C2H2
       (b) Sulphur molecule, S8
       (c) Phosphorus molecule, P4 (Atomic mass of phosphorus = 31)
       (d) Hydrochloric acid, HCl
       (e) Nitric acid, HNO3

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


The molar mass of the following: [Unit is ‘g’]
(a) Ethyne, C2H2= 2 × 12 + 2 × 1 = 24 + 2 = 26 g
(b) Sulphur molecule, S8= 8 × 32 = 256 g
(c) Phosphorus molecule, P4= 4 × 31 = 124 g
(d) Hydrochloric acid, HCl = 1 × 1 + 1 × 35.5
= 1 + 35.5 = 36.5 g
(e) Nitric acid, HNO3= 1 × 1 + 1 × 14 + 3 × 16
= 1 + 14 + 48 = 63 g



Thanks